# Persistent Cost of Lipschitz Maps

### v1

@@ Abstract
A 1-Lipschitz map between compact metric spaces $f:X\to Y$ induces a homomorphism on the level of persistent d-homology of the associated Vietoris-Rips simplicial filtrations. This homomorphism naturally bounds the Bottleneck distance between said modules in terms of the maximum size of its kernel and cokernel persistence modules, a stable invariant which we refer to as the persistent cost associated to $f$. We provide metric upper bounds on the persistent cost of a 1-Lipschitz map and revisit its stability.
@@ Introduction
Persistent homology is a fundamental tool in the emerging field of topological data analysis (TDA) that extracts a vast amount of metric information by encoding homological features across scales. The richness of this output coupled with its explainability has yielded significant insights across a spectrum of disciplines, including biology and time series analysis, among others, c.f.[otter2017roadmap, perea2015sliding, gardner2022toroidal], and been integrated into learning pipelines, c.f., [hensel2021survey, hofer2017deep].
We outline the standard pipeline for topological data analysis. Given a metric space $X$, the Vietoris–Rips filtration $(\operatorname{VR}_{s}(X))_{s\geq 0}$ is constructed by including simplices whose vertex set has diameter at at most $s\geq 0$. As the parameter increases, the complex evolves from a discrete set to the full power set of $X$, encoding geometric information at multiple scales. Applying the homology functor in a fixed dimension $d$ yields a persistence module: a sequence of vector spaces connected by linear maps that track the birth, persistence and death of homological features.
While the TDA community has primarily focused on analyzing individual point clouds, the functorial nature of these constructions naturally leads us to consider mappings between datasets. Indeed, we advocate for a treatment of maps between spaces that parallels that of the spaces themselves, enabling us to address the geometry of functions from a persistent homology perspective. In this context we shall consider the natural case of non-expanding, i.e., 1-Lipschitz, maps $f:X\to Y$ between Vietoris-Rips filtrations of finite metric spaces which yields a homomorphism of point-wise finite dimensional persistence modules,
$f_{*}:\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}}(X))\to\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}}(Y)).$
@@
This persistent cost bounds the interleaving distance between the modules associated to $X$ and $Y$, by the Induced Matching Theorem of U. Bauer and M. Lesnick in [bauer2015induced] (see also [bauer2020persistence]).
Building on this result, we contribute a metric upper bound for the persistent cost of non-expanding map $f:X\to Y$, showing that $C(f)$ is bounded above by its distortion and the Hausdorff distance between its image and the codomain, $f(X)\subseteq Y$. Specifically, we establish the second inequality in the following sequence.
$d_{I}\left(\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}}(X)),\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}}(Y))\right)\penalty 10000\ \leq\penalty 10000\ C(f)\penalty 10000\ \leq\penalty 10000\ \operatorname{dist}(f)+2\cdot d_{H}(f(X),Y)$ (2)
@@
A prior concern is the stability of kernel and cokernel persistence modules which was addressed in [cohen2009persistent]. We provide a self-contained exposition of this fact that was claimed without proof. In doing so, we introduce a Gromov-Hausdorff type distance between maps $f:X\to Y$, $f^{\prime}:X^{\prime}\to Y^{\prime}$, of compact metric spaces that may be of independent interest.
Our results are developed within the framework of point-wise finite dimensional persistence modules and finite metric spaces, and naturally extended to q-tame modules that arrise from compact metric spaces.
The paper is organized as follows. A first section provides a self-contained exposition of the algebraic notions regarding persistence theory. We then move on to the geometric of spaces, maps and distances. The last two subsection are focused on our contributed results on metric upper bounds for the cost, Theorem 2.11 and stability, Theorem 2.16.
@@ 1 Algebraic Persistence
The main references for this section are the comprehensive book by S. Oudot in [oudot2015persistence] and the papers on the Induced Matching Theorem [bauer2015induced, bauer2020persistence] by U. Bauer and M. Lesnick.
@@ 1.1 Persistence Modules
where $\mathrm{Vect}_{\mathbb{F}}$ denotes the category of vector spaces over a field $\mathbb{F}$ with linear transformations as homomorphisms.
This, a persistence module is more than just a one-parameter family of vector spaces; as a representation of a poset, it carries “inner persistence” or “transition” maps
$\varphi_{s<t}:V_{s}\xrightarrow{\penalty 10000\ \penalty 10000\ }V_{t},$
for each pair $s<t\in\mathbb{R}$, which satisfy
@@
As a first and computationally relevant approach to the subject, the reader may consider the case of finite-dimensional spaces indexed over a finite set $0=\epsilon_{0}<\epsilon_{2}<\dots<\epsilon_{r}$, giving rise to
$V_{\epsilon_{0}}\to V_{\epsilon_{1}}\to\cdots\to V_{\epsilon_{r}}.$ (4)
The key idea of persistence is that the dimension of $V_{s}$ may be preserved (via an isomorphism) in $V_{t}$ for $t>s$.
However, indexing over the real numbers has several advantages. It is less cumbersome, as it avoids the problem of refining different indexing sets. It is also more general, since any finitely indexed persistence module can be easily extended to a $\mathbb{R}$-indexed one.
We will first develop the theory within the setting of point-wise finite dimensional (p.f.d.) modules indexed over $\mathbb{R}$, as in definition (3). A fundamental result states that such persistence modules admit a unique decomposition into interval modules [crawley2015decomposition, oudot2015persistence]. Specifically,
$V\cong\bigoplus_{I\in\mathcal{I}}\mathbb{F}^{I},$ (5)
where $\mathcal{I}$ is a collection of intervals $I\subseteq\mathbb{R}$ each defining an “interval module” as follows:
$(\mathbb{F}^{I})_{(t)}=\begin{cases}\mathbb{F},&t\in I,\\ 0,&\text{otherwise},\end{cases}$
with inner persistence maps:
$\varphi_{s<t}=\begin{cases}Id,&s,t\in I,\\ 0,&\text{otherwise}.\end{cases}$
The field $\mathbb{F}$ will play no role in the discussion, usually being $\mathbb{Z}_{p}$ for p prime, $\mathbb{Q}$ or $\mathbb{R}$ in practice.
The multiset of intervals $\mathcal{I}$ is called the persistence diagram (or barcode). Each interval $\mathcal{I}$ is represented buy its endpoints $(b_{i},d_{i})\in\bar{\mathbb{R}}^{2}$. A barcode represents each interval as a horizontal line from $b$ to $d$, as shown in Figure 1.
bbdd
12345tt
Figure 1: A persistence diagram and barcode representation of the same decomposition.
In order to define a distance between persistence diagrams $D_{1},D_{2}$, we may begin by setting a partial matching between their persistent features. If a pair $(b_{1},d_{1})\in D_{1}$ and $(b_{2},d_{2})\in D_{2}$ are matched, we may ponder their dissimilarity by the quantity $(b_{1}-b_{2},d_{1}-d_{2})||_{\infty}$. Unmatched elements $(b,d)\in D_{i}$ are virtually paired to an ephemerous feature $(\frac{b+d}{2},\frac{b+d}{2})$, i.e., to an element listed in neither diagram, thus weighting in as $\frac{(d-b)}{2}$. In other words, for a partial matching $\Theta$ between diagrams $D_{1},D_{2}$ we have defined its dissimilarity as
$diss(\Theta_{D_{1},D_{2}})=\max\Bigg\{\underset{(b_{1},d_{1})\sim_{\Theta}(b_{2},d_{2})}{\sup\quad\quad}||(b_{1}-b_{2},d_{1}-d_{2})||_{\infty}\penalty 10000\ \penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \underset{(b,d)\notin\pi_{i}(\Theta)}{\sup\quad\quad}\frac{(d-b)}{2},\penalty 10000\ \Bigg\}$
Definition 1.1.
The Bottleneck distance between persistence diagrams $D_{1},D_{2}$ is defined as the infimum dissimilarity over all partial matchings, i.e.,
$d_{Bottl}(D_{1},D_{2})=\underset{\theta}{\inf}\penalty 10000\ diss(\theta_{D_{1},D_{2}})$ (6)
Notice that any interval decomposable persistence module is at zero distance of a module constructed using only intervals of the form $[a,b)$. We will comment on this in Section 1.4
@@ 1.2 Homomorphisms and Interleavings
A homomorphism $F:V_{\bullet}\to W_{\bullet}$ between persistence modules is a natural transformations between $V_{\bullet}$ and $W_{\bullet}$ regarded as functions. In other words, a homomorphism is a collection of linear maps $F_{\bullet}$, at each scale, commuting with inner maps of each module, so that for $s<t$ we have:
@@
In practice, it is $\epsilon\geq 0$, though it is not necessary.
We may compare modules by means of interleavings which are defined as follows.
@@
Whenever clear from context we omit indexing homomorphisms and naming inner persistence maps of each module.
@@
$d_{I}(V_{{}_{\bullet}},W_{{}_{\bullet}})=\inf\{\epsilon:\text{there is an $\epsilon$-interleaving between $V_{\bullet}$ and $W_{\bullet}$}\}$
The interleaving distance agrees with the Bottleneck distance between associated diagrams, in the interval decomposable case, i.e.,
$d_{Bottl}(diag(V_{\bullet}),diag(W_{\bullet})=d_{I}(V_{\bullet},W_{\bullet}).$
This result is known as the algebraic isometry theorem, see [lesnick2015theory, oudot2015persistence].
1.3 Kernel, cokernel and the Induced Matching Theorem
Given a homomorphism $F:V_{\bullet}\to W_{\bullet}$, we can define persistence modules for the point-wise kernel, cokernel, and image. Standard diagram chasing shows that these form an exact sequence of persistence modules.
@@ 1.3 Kernel, cokernel and the Induced Matching Theorem
The following result describes how a homomorphism induces a matching between the interval decompositions of $V_{\bullet}$ and $W_{\bullet}$.
Theorem 1.5 (Induced Matching Theorem, [bauer2015induced, bauer2020persistence]).
Let $V_{\bullet},W_{\bullet}$ be p.f.d. persistence modules and $F_{\bullet}:V_{\bullet}\to W_{\bullet}$ a homomorphism. Then there exists an induced matching between intervals in their respective interval decompositions satisfying the following:
  1. 1.
    a matched pair of bars $[a,b),[c,d)$ in the respective decompositions of $V_{\bullet}$ and $W_{\bullet}$, satisfies $a\leq c<b\leq d$;
  2. 2.
    the unmatched intervals in $V_{\bullet}$, as well as the non-overlapping segment $[a,c)$ of a pair of matched bars, correspond both to intervals in $\operatorname{Ker}_{\bullet}$/
  3. 3.
    analogously, the unmatched intervals in $W_{\bullet}$, as well as the non-overlapping segment $[b,d)$ of a pair of matched bars, correspond both to intervals in $\operatorname{Coker}_{\bullet}$;
  4. 4.
    the middle portion $[c,b)$ in common to a matched pair of bars.
1.4 The right setup
We conclude this algebraic section by reviewing the framework of q-tame persistence modules, following [chazal2014observable] and [oudot2015persistence].
A q-tame persistence module over $\mathbb{R}$ is a slightly more general notion of persistence module where the point-wise finite dimensional hypothesis is replaced by the weaker requirement that every inner persistence map $\varphi_{s<t}$ be of finite rank.
A p.f.d. module can be retrieved from a q-tame one, for example one could consider $\hat{V}_{t_{0}}:=\lim_{s\to{t_{0}}^{+}}\operatorname{Im}(\varphi_{t_{0}<s})$. In fact there is a quotient category, the observable category, retrieved from the stabilization of persistence modules up to direct sum with the “ephemeral” subcategory given by those persistence modules all of whose inner homomorphisms are trivial, i.e.,
$\mathbf{Obs}:=\mathbf{Pers}/\mathbf{Eph},$
where $\mathbf{Pers}:=[(\mathbb{R},\leq),\mathbf{Vect}_{\mathbb{F}}]$ as before. Ephemerous modules correspond to “still-born” persistence features of the form $[a,a]$.
Every q-tame persistence module is isomorphic to a p.f.d. one in $\mathbf{Obs}$, particularly to a unique interval-decomposable one with intervals of chosen decoration, i.e., type of strict or relaxed condition on endpoints: $[a,b],(a,b),[a,b)$ or $(a,b]$. The barcode associated to a q-tame module is understood to be that of an interval decomposition of its associated p.f.d. observable.
The homology of the Vietoris-Rips filtration over compact metric may not yield a p.f.d. persistence module but a q-tame one nonetheless.
Throughout the text "distance" is understood in the general sense as a symmetric non-negative extended real function on pairs for which the triangular inequality holds. However, it may only be a pseudo-distance in the sense that there is a degenerate non-unique set of points at zero distance. Considering the observable category solves the issue, as the pseudo-distance becomes a true distance once zero interleaved p.f.d. persistence modules become isomorphic in $\mathbf{Obs}$, see [lesnick2015theory]. In particular, this makes the already introduced graphic representation as persistence diagrams or barcodes a faithful one, as in previous Figure 1.
2 Persistent cost
@@
A general reference for this section is Oudot’s book [oudot2015persistence] with regards to persistence theory and the references therein. For metric arguments we refer to [burago2001metricgeometry].
2.1 Vietoris-Rips filtration and maps
The Vietoris-Rips construction is a standard procedure by which a filtration of simplicial complexes is produced from an actual dataset, presented as a finite metric space. To be precise, for each $\epsilon\geq 0$ define:
$\mathrm{VR}_{\epsilon}(X)=\{\sigma\subseteq X\mid d(x,y)\leq\epsilon,\,\forall x,y\in\sigma\}\subseteq\mathcal{P}(X),$
i.e., a collection of subsets of $X$. Applying the d-dimensional homology functor $H_{d}(-,\mathbb{F)}$, with coefficients in a field $\mathbb{F}$, yields a persistence module $V_{\bullet}:=\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))$ with inner persistence maps $\varphi_{s<t}:=(\iota_{s<t})_{*}$.
@@ 2.1 Vietoris-Rips filtration and maps
$d_{Y}(h(x),h(x^{\prime})\leq d_{X}(x,x^{\prime})+\epsilon.$
The Vietoris-Rips construction is particularly well suited for our metric endeavors, as follows from the following.
@@
Let $h:X\to Y$ be an $\epsilon$-quasi $1$-Lipschitz map. Then $h$ induces an $\epsilon$-degree persistence homomorphism: $h_{*}:H_{d}(\operatorname{VR}_{{}_{\bullet}}(X))\to H_{d}(\operatorname{VR}_{{}_{\bullet}+\epsilon}(Y)).$
@@
$\hat{h}:\operatorname{VR}_{{}_{\bullet}}(X)\to\operatorname{VR}_{{}_{\bullet},+\epsilon}(Y),\quad[x_{1}\dots x_{r}]\mapsto[h(x_{1})\dots h(x_{r})],$
since a set $\sigma\subseteq X$ of diameter $\ell$ maps to a set $h(\sigma)\subseteq Y$ of diameter at most $\ell+\epsilon$. This commutes with inclusions and thus gives a morphism between filtrations. The homology functor yields the desired persistence homomorphism.
For the second claim, suppose $d_{Y}(h_{1}(x),h_{2}(x))\leq\delta$ for all $x\in X$. For any simplex $\sigma\in\operatorname{VR}_{{}_{\bullet}}(X)$ of diameter $\ell$, consider the union:
@@
$\operatorname{VR}_{{}_{\bullet}}(X)\to\operatorname{VR}_{{}_{\bullet}+\epsilon+\delta}(Y),$
@@
$f_{*}:H_{d}(\operatorname{VR}_{{}_{\bullet}}(X))\to H_{d}(\operatorname{VR}_{{}_{\bullet}}(Y)).$
@@
A source of $\epsilon$-quasi $1$-Lipschitz maps with non-trivial $\epsilon>0$ is given by the choice of functions subordinated to a given correspondence between metric spaces.
@@
A correspondence between two sets is a relation between two subsets, $\mathcal{C}\subseteq X\times Y$, that in fact relates every element of the factors, i.e., such that $\pi_{X}(\mathcal{C})=X$, $\pi_{Y}(\mathcal{C})=Y$.
When these spaces $x,y$, are metric we have a notion of distortion given by the supremum of the difference between the distances of related pairs, i.e.,
@@
Given metric spaces $X$, $Y$, consider all possible correspondences and take half the infimum over their distortions to be the Gromov-Hausdorff distance between them,
$d_{GH}(X,Y)=\frac{1}{2}\,\underset{\mathcal{C}}{\inf}\{\!\penalty 10000\ dist(\mathcal{C})\!\penalty 10000\ \}.$
The alternative, albeit original, definition is to consider the infimum over all metric spaces $Z$ and isometric embeddings $f:X\to Z$, $g:Y\to Z$, of the quantity $d_{\mathrm{H}}^{Z}$ given by the Hausdorff distance in $Z$ of the respective images from $X,Y$.
$d_{\mathrm{GH}}(X,Y)=\inf_{Z,f,g}d_{\mathrm{H}}^{Z}(f(X),g(Y)).$
The Gromov-Hausdorff distance is a distance when we consider compact subsets up to isometry.
@@
Theorem 2.7 (Stability [cohen2005stability]).
@@
$d_{I}(\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X),\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\leq 2\,d_{GH}(X,Y).$
This key result was first proven by Cohen-Steiner, Edelsbrunner and Harer, as duly cited, c.f. [oudot2015persistence]. It provides a central feature of persistent homology which is its stability, namely, the fact that the considered invariants are continuous on the inputs.
@@
The co-restriction of a given map between metric spaces, $f|:X\to f(X)$, being automatically surjective is already a correspondence. Its classical distortion can be interpreted as the absolute error between the distances of inputs and outputs under $f$,
@@
An inclusion of compact metric spaces $A\subseteq Y$ can be enlarged to a correspondence $C_{\iota}$ by adding all the pairs $(a,y)\in A\times Y$ for $a=argmin\penalty 10000\ d(.,y)$. Its distortion is (sharply) bounded by twice the Hausdorff distance from $A$ to $Y$, i.e.,
@@
Given that a composition of correspondences is again a correspondence, the previous two examples show that a any function $f:X\to Y$ between compact metric spaces can be enlarged, as a relation, to induce a correspondence between $X$ and $Y$, so that:
$\mathcal{C}_{f}=\{(x,y)\in X\times Y\penalty 10000\ :\penalty 10000\ x=\operatorname{argmin}\penalty 10000\ d(f(.),y)\},$
so that the $x$ that pair to an $y\in Y$ are the arguments that minimize distance to $y$ via $f$.
Moreover, as a composition of the previous examples its distortion is bounded as follows:
@@
2.2 Bounds on the Persistent Cost
We have all the necessary ingredient to take on the subject of this work which is the persistent cost associated to a non-expanding map between compact metric spaces. We shall provide natural upper bounds that control it and stability guarantees.

### v2

# arXiv:2511.07674v2[math.AT] 02 Mar 2026

@@ Abstract
A $1$-Lipschitz map between compact metric spaces $f\colon X\to Y$ induces a homomorphism of persistence modules on degree-$d$ Vietoris–Rips persistent homology. We define the persistent cost of $f$ from this induced homomorphism by quantifying the persistence carried by its kernel and cokernel modules.
We prove that the persistent cost controls the interleaving distance between the degree-$d$ Vietoris–Rips persistent homology modules of $X$ and $Y$. Moreover, we obtain an explicit upper bound for the persistent cost in purely metric terms. Finally, we give a self-contained proof of the stability of the persistent cost introducing a Gromov-Hausdorff type distance for maps between compact metric spaces.
@@ Introduction
Persistent homology is a fundamental tool in the emerging field of topological data analysis (TDA) that extracts a vast amount of metric information by encoding homological features across scales. The richness of this output coupled with its explainability has yielded significant insights across a spectrum of disciplines, including biology and time series analysis, among others, c.f. [OPT+17, PH15, GHP+22], and been integrated into learning pipelines, c.f. [HMR21, HKN+17].
Given a metric space $X$, the Vietoris–Rips filtration $(\operatorname{VR}_{s}(X))_{s\geq 0}$ is constructed by including simplices whose vertex set has diameter at most $s\geq 0$. As the parameter increases, the complex evolves from a discrete set to the full power set of $X$, encoding geometric information at multiple scales. Applying the homology functor in a fixed dimension $d$ yields a persistence module: a sequence of vector spaces connected by linear maps that track the birth, persistence and death of homological features.
While the TDA community has primarily focused on analyzing individual point clouds, the functorial nature of these constructions naturally leads us to consider maps between datasets. Indeed, we advocate for a treatment of maps between spaces analogous to that of the spaces themselves, enabling us to address the geometry of functions from a persistent homology perspective. In this context we shall consider the natural case of non-expanding, i.e., 1-Lipschitz, maps $f:X\to Y$ which induce morphisms between their Vietoris–Rips filtrations and hence homomorphisms of persistence modules,
$f_{*}:\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))\to\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y)).$
@@
The persistent cost bounds the interleaving distance between the modules associated to $X$ and $Y$, by a technical improvement to the Induced Matching Theorem of U. Bauer and M. Lesnick in [BL15], see Theorem 1.5.
On the other hand, we can explicit a metric upper bound for the persistent cost of a non-expanding map $f:X\to Y$, in terms of its distortion and Hausdorff distance between its image subset and codomain, $f(X)\subseteq Y$. Altogether, we establish the following sequence of inequalities:
$d_{I}\left(\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X)),\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\right)\penalty 10000\ \leq\penalty 10000\ C(f)\penalty 10000\ \leq\penalty 10000\ \operatorname{dist}(f)+2\cdot d_{H}(f(X),Y)$ (2)
We show that this bounds are sharp.
@@
A related issue concerns the stability of kernel and cokernel persistence modules, which was claimed without proof in [CEH+09] as a corollary of the machinery developed in [CEH05]. We provide here a self-contained exposition of this fact introducing a Gromov-Hausdorff type distance for maps $f:X\to Y$, $f^{\prime}:X^{\prime}\to Y^{\prime}$, between compact metric spaces, that may be of independent interest.
The paper is organized as follows. A first section provides a minimal exposition of the algebraic notions that allow us to define the persistent cost and prove the first inequality above, Theorem 1.5. We then move on to the geometry of spaces and maps, in order to establish the metric upper bound on Theorem 2.11 and stability on Theorem 2.16.
@@ 1 Algebraic Persistence
The main references for this section are the comprehensive book by S. Oudot in [OUD15] and the papers on the Induced Matching Theorem [BL15, BL20] by U. Bauer and M. Lesnick.
@@ 1.1 Persistence Modules
where $\mathrm{Vect}_{\mathbb{F}}$ denotes the category of vector spaces over a field $\mathbb{F}$ with linear transformations as homomorphisms. Thus, a persistence module is more than just a one-parameter family of vector spaces; as a representation of a poset, it carries “inner persistence” or “transition” maps $\varphi_{s<t}:V_{s}\xrightarrow{\penalty 10000\ \penalty 10000\ }V_{t}$, for each pair $s<t\in\mathbb{R}$, which satisfy:
@@
Additional hypotheses are often imposed on persistence modules to ensure that their representation category has satisfactory structural properties. A possible choice is to restrict to pointwise finite-dimensional (p.f.d.) modules, which admit an interval decomposition due to the work in [CRA15]. A broader condition is $q$-tameness, requiring that each internal structure map $\varphi_{s<t}$ has finite rank.
[…]
$V^{\epsilon}_{r}=\mathrm{im}\,\varphi_{r-\epsilon<r}$ (4)
has finite dimension. In particular, $V^{\epsilon}_{\bullet}$ is a p.f.d. submodule of $V_{\bullet}$.
Furthermore, there is a natural $\epsilon$-interleaving between a module $V_{\bullet}$ and its $\epsilon$-approximation $V^{\epsilon}_{\bullet}$, the pair of maps being given by the inclusion $V^{\epsilon}_{\bullet}\subseteq V_{\bullet}$ and inner persistence from $\varphi_{\bullet<\bullet+\epsilon}:V_{\bullet}\to V^{\epsilon}_{\bullet+\epsilon}$.
@@ 1.2 Homomorphisms and Interleavings
A homomorphism $F:V_{\bullet}\to W_{\bullet}$ between persistence modules is a natural transformation between $V_{\bullet}$ and $W_{\bullet}$ regarded as functors.
In other words, a homomorphism is a collection of linear maps $F_{\bullet}$, at each scale, commuting with inner maps of each module, so that for $s<t$ we have the diagram to the side.
@@
Whenever it is clear from context we omit indexing homomorphisms and naming inner persistence maps of each module.
@@
$d_{I}(V_{\bullet},W_{\bullet})=\inf\{\epsilon\geq 0:\text{there is an $\epsilon$-interleaving between $V_{\bullet}$ and $W_{\bullet}$}\}$
Though referred as “distance”, it is only an extended pseudo-distance, i.e., it may attain the value $\infty$ and assign zero distance to different modules.
The interleaving distance agrees with the Bottleneck distance between associated diagrams, in the interval decomposable case where the latter is defined. This result is known as the algebraic isometry theorem, see [LES15, OUD15].
1.3 Persistent Cost
Associated to a homomorphism $F:V_{\bullet}\to W_{\bullet}$, we can define persistence modules given as the point-wise kernel, cokernel, and image spaces, with persistence maps naturally induced from those of $V_{\bullet}$ and $W_{\bullet}$. In particular, we retrieve an exact sequence of persistence modules given by:
@@ 1.3 Persistent Cost
The size of the kernel and cokernel encode how far $F$ is from being a persistence isomorphism, motivating the following.
Definition 1.4.
The persistent cost associated to a homomorphism $F:V_{\bullet}\to W_{\bullet}$ of persistence modules is the larger of the interleaving distances of the kernel and the cokernel to the trivial persistence module $\Delta$,
$C(F)=\max\left\{d_{I}(\operatorname{Ker}(F),\Delta),\ d_{I}(\operatorname{Coker}(F),\Delta)\right\}.$
We may now state our first contribution.
Theorem 1.5.
Let $F:V_{\bullet}\to W_{\bullet}$ be a homomorphism between q-tame persistence modules, then its persistent cost bounds the interleaving distance between $V_{\bullet}$ and $W_{\bullet}$, i.e.,
$d_{I}(V_{\bullet},W_{\bullet})\leq C(F).$
Proof. If the modules are p.f.d. the statement follows as a corollary to the Induced Matching Theorem of U. Bauer and M. Lesnick in [BL15].
Recall the auxiliary definition of $V^{\epsilon}_{\bullet}$ in (4) and observe that a persistence homomorphism $F$ corestricts well to:
$F^{\epsilon}:=F|:V^{\epsilon}_{\bullet}\to W^{\epsilon}_{\bullet}.$
In particular from [BL15] we have:
$d_{I}(V^{\epsilon}_{\bullet},W^{\epsilon}_{\bullet})\leq C(F^{\epsilon}).$
We need only check that:
$d_{I}(\operatorname{Ker}(F),\operatorname{Ker}(F^{\epsilon}))=\epsilon\quad,\quad d_{I}(\operatorname{Coker}(F),\operatorname{Coker}(F^{\epsilon})=\epsilon,$
where the pair of maps that give the desired interleavings are induced from those in (4) between each module $V_{\bullet}$, $W_{\bullet}$ and its corresponding $\epsilon$-approximation $V^{\epsilon}_{\bullet}$, $W^{\epsilon}_{\bullet}$.
A straight-forward triangular inequality gives:
$d_{I}(V_{\bullet},W_{\bullet})\leq 2\epsilon+C(F^{\epsilon})\leq 4\epsilon+C(F).$
Given that $\epsilon>0$ can be set up independently, the theorem follows. $\square$
Notice that this bound is sharp, a simple example given by the null homomorphism from the trivial module into any other.
2 On maps
@@
A general reference for this section is the book by Oudot [OUD15] with regards to persistence theory. Our metric arguments elaborate on classical metric geometry, for which we refer to [BBI01].
2.1 Quasi Lipschitz maps
@@
$d_{Y}\big(h(x),h(x^{\prime})\big)\leq d_{X}\big(x,x^{\prime}\big)+\epsilon.$
The Vietoris-Rips construction is particularly well suited for our metric endeavors, as the following result shows.
@@
Let $h:X\to Y$ be an $\epsilon$-quasi $1$-Lipschitz map. Then $h$ induces an $\epsilon$-degree persistence homomorphism: $h_{*}:H_{d}(\operatorname{VR}_{\bullet}(X))\to H_{d}(\operatorname{VR}_{\bullet+\epsilon}(Y)).$
@@
$\hat{h}:\operatorname{VR}_{\bullet}(X)\to\operatorname{VR}_{\bullet+\epsilon}(Y),\quad[x_{0}\dots x_{r}]\mapsto[h(x_{0})\dots h(x_{r})],$
since a set $\sigma\subseteq X$ of diameter $\ell$ maps to a set $h(\sigma)\subseteq Y$ of diameter at most $\ell+\epsilon$. This commutes with inclusions and thus gives a homomorphism between filtrations. The homology functor yields the desired persistence homomorphism.
For the second claim, suppose $d_{Y}(h_{1}(x),h_{2}(x))\leq\delta$ for all $x\in X$. For any simplex $\sigma\in\operatorname{VR}_{\bullet}(X)$ of diameter $\ell$, consider the union:
@@
$\operatorname{VR}_{\bullet}(X)\to\operatorname{VR}_{\bullet+\epsilon+\delta}(Y),$
@@
$f_{*}:H_{d}(\operatorname{VR}_{\bullet}(X))\to H_{d}(\operatorname{VR}_{\bullet}(Y)).$
@@
A correspondence is a relation between two sets, $\mathcal{C}\subseteq X\times Y$, that relates every element of the factors, i.e., such that $\pi_{X}(\mathcal{C})=X$, $\pi_{Y}(\mathcal{C})=Y$.
When these spaces $X,Y$, are metric one can define the distortion associated to a correspondence as the supremum of the difference between the distances of related pairs, i.e.,
@@
Given metric spaces $X$, $Y$, the Gromov-Hausdorff distance between them is defined by,
$d_{GH}(X,Y)=\frac{1}{2}\,\underset{\mathcal{C}}{\inf}\{\!\penalty 10000\ dist(\mathcal{C})\!\penalty 10000\ \},$
where the infimum is taken over all possible correspondences.
The alternative, albeit original, definition is to consider the infimum of the Hausdorff distance between isometric copies of $X,Y$:
$d_{\mathrm{GH}}(X,Y)=\inf_{Z,f,g}d_{\mathrm{H}}^{Z}(f(X),g(Y)),$
where the infimum ranges over isometric embeddings $f:X\to Z$, $g:Y\to Z$ into a common ambient space $Z$.
The Gromov-Hausdorff distance is a distance when we consider compact metric spaces up to isometry.
@@
Theorem 2.7 (Stability [CEH05]).
@@
$d_{I}\big(\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X)),\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\big)\leq 2\,d_{GH}(X,Y).$
This key result provides a central feature of persistent homology which is its stability, namely, the fact that the considered invariants are continuous on the inputs.
@@
The co-restriction of a given map between metric spaces, $f|:X\to f(X)$, being automatically surjective is already a correspondence. Its distortion can be interpreted as the absolute error between the distances of inputs and outputs under $f$,
@@
An inclusion of compact metric spaces $A\subseteq Y$ can be enlarged to a correspondence $C_{\iota}$ by adding all the pairs $(a,y)\in A\times Y$ for $a=argmin\penalty 10000\ d(.,y)$. Its distortion is (sharply) bounded by twice the Hausdorff distance between $A$ and $Y$, i.e.,
@@
Given that a composition of correspondences is a correspondence, the previous two examples show that a any function $f:X\to Y$ between compact metric spaces can be enlarged as a relation to induce a correspondence between $X$ and $Y$:
$\mathcal{C}_{f}=\{(x,y)\in X\times Y\penalty 10000\ :\penalty 10000\ x=\operatorname{argmin}\penalty 10000\ d(f(.),y)\}.$
Moreover, as a composition its distortion is bounded as follows:
@@
2.2 A metric upper bound
We now have the necessary framework to address the main topic of this work which is the persistent cost associated to a non-expanding map between compact metric spaces. We shall provide natural upper bounds and stability guarantees.